OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. 1. cot x sec4x = cot x + 2 tan x + tan3x

zepdrix (zepdrix):

Are those exponents or coefficients? \(\large \cot x\sec^4x\) or \(\large \cot x\sec4x\) ?

OpenStudy (anonymous):

exponents

zepdrix (zepdrix):

Oh ok, now this problem makes more sense :)

zepdrix (zepdrix):

Here is an important identity we'll want to use, \[\large \color{green}{\sec^2x=1+\tan^2x}\] So let's start here,\[\large \cot x\left(\sec^4x\right) \qquad=\qquad \cot x \left(\color{green}{\sec^2x}\right)^2\] Understand what I did with the exponent?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

did u factor it

zepdrix (zepdrix):

Rule of exponents: \(\large (x^a)^b =x^{ab}\) We're using this rule in reverse, \(\large x^4\qquad=\qquad x^{2\cdot2}\qquad=\qquad (x^2)^2\)

OpenStudy (anonymous):

ok i get it now

zepdrix (zepdrix):

I colored the inside portion green, see how we'll apply the identity?

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\large \cot x \left(\color{green}{\sec^2x}\right)^2\qquad=\qquad \cot x \left(\color{green}{1+\tan^2x}\right)^2\] From here, expand the outer square, the black one.

zepdrix (zepdrix):

In case you're confused about what I'm asking, \[\large \cot x \left(\color{green}{1+\tan^2x}\right)^2\qquad=\qquad \cot x \left(\color{green}{1+\tan^2x}\right)\left(\color{green}{1+\tan^2x}\right)\] The next step is to multiply out the brackets.

OpenStudy (anonymous):

ok